OpenStudy (anonymous):

segment ab and segment cd are chords of circle p intersecting at x. if ax=7.5, bx=3.2, cd=11 and cx>dx find cx.

jimthompson5910 (jim_thompson5910):

let y be the length of cx so cx + dx = cd y + dx = 11 dx = 11 - y

jimthompson5910 (jim_thompson5910):

now use the intersecting chords theorem to get ax*bx = cx*dx 7.5*3.2 = y*(11-y) solve for y